OpenStudy (anonymous):

Given that h(x) = 4x+5 and (h-j)(x) = x-2, find j(x).

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

I think I have an idea, but I don't think it is correct

OpenStudy (anonymous):

Would I do 4(x-2)+5

OpenStudy (mathstudent55):

h(x) = 4x+5 and (h-j)(x) = x-2, find j(x) (h-j)(x) = h(x) - j(x) x - 2 = 4x + 5 - j(x) Solve for j(x)

OpenStudy (anonymous):

Ok. Would that be -7/3 ?

OpenStudy (mathstudent55):

x - 2 = 4x + 5 - j(x) Add j(x) to both sides: j(x) + x - 2 = 4x + 5 Subtract x from both sides & add 2 to both sides: j(x) = 3x + 7

OpenStudy (anonymous):

Oh ok.. And then once I plug them in together I should get x-2 right?

OpenStudy (mathstudent55):

Let's check. We must have h(x) - j(x) = x - 2 h(x) = 4x + 5 j(x) = 3x + 7 (h-j)(x) = h(x) - j(x) = 4x + 5 - (3x + 7) = 4x + 5 - 3x - 7 = x - 2

OpenStudy (anonymous):

Thank you so much! I thought it was something along the likes of that, but I just wanted to double check.

OpenStudy (mathstudent55):

Yes, but the answer is just j(x) = x - 2 I just plugged it in to check if j(x) = x - 2 was the correct answer.

OpenStudy (mathstudent55):

You're welcome.